OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely. Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3

OpenStudy (anonymous):

for \(S_k\) replace every \(n\) by \(k\)

OpenStudy (anonymous):

So.... 60.. i am not understanding how u replace n by k

OpenStudy (anonymous):

the answer is not a number write down exactly what you see written above, but instead of putting an \(n\) put a \(k\)

OpenStudy (anonymous):

the left hand side of the equal sign will be \[1 \times 2 + 2 \times 3 + 3 \times 4 + . . . + k(k + 1) \]

OpenStudy (anonymous):

so unless this is it im still confused..Sk: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + k(k + 1) = [k(k + 1)(k + 2)]/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

then for the next one, replace all the \(n\)'s by \(k+1\) make sure to use parentheses and the distributive law when you simplify

OpenStudy (anonymous):

ok thank you